OpenStudy (anonymous):

what is the standard deviation of 1 2 3 6 8?

OpenStudy (anonymous):

Can you find the mean, first?

OpenStudy (anonymous):

first find the mean: \[ \mu = \frac{1+2+3+6+8}{5} \]

OpenStudy (anonymous):

is the answer 4?

OpenStudy (anonymous):

Then find std dev: \[ \sigma = \sqrt{\frac{(1-\mu)^2+(2-\mu)^2+(3-\mu)^2+(6-\mu)^2+(8-\mu)^2}{5}} \]

OpenStudy (anonymous):

No, 4 would be \(\mu\), the mean

OpenStudy (anonymous):

Then plug that into the equation I gave you for \(\sigma\).

OpenStudy (anonymous):

is it 2.607680962

OpenStudy (anonymous):

thank you! so it would be 2.9

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

In a sample of 800 people, 32 people said their favorite fruit was an apple. Determine the empirical probalility that a person's favorite fruit is an apple.

OpenStudy (anonymous):

can you help me with that?